OpenStudy (anonymous):

so i just joined my calculus class about 4 days late and i'm already behind and i'd love if someone could help me run through how to do limit problems what is the limit when x approaches 0 at sin 1/2x all over x

jimthompson5910 (jim_thompson5910):

This is the limit problem? \[\Large \lim_{x \to 0} \frac{\sin\left(\frac{1}{2}x\right)}{x}\]

OpenStudy (dinamix):

the solution is 1/2

OpenStudy (dinamix):